A politician wants to know his chances of winning in the next election. He has tasked a researcher to interview a random sample of his contituents about their perceptions of his leadership. Previous polls have revealed that approximately 62.5% are in favor of his leadership. The politician has adopted the 90% confidence level and of the population value by at most 0.02. How many respondents must be interviewed?


1
Expert's answer
2021-05-13T11:56:33-0400

"p\:=62.5\%=\:0.625"

"Margin\:of\:error,E=0.02"

"The\:critical\:value\:for\:the\:significance\:level\:\alpha \:=\:0.1\:is\:z_c\:=\:1.645,"

"which\:is\:obtained\:by\:looking\:into\:a\:standard\:normal\:probability\:table."

"Sample size, n\u200b\\:\\:\\ge \u200b\\:\\:p\\left(1\u2212p\\right)\\left(\\:\\frac{Zc}{E}\\:\u200b\\:\\right)^2"

"=0.625\\left(1\u22120.625\\right)\\left(\\:\\frac{1.645}{0.02}\\:\u200b\\:\\right)^2\\:"

"=1585.28\:\approx 1586"

"Required\:sample\:size,\:n=1586"